A photon having E = 37.7 keV energy scatters from
Answer #1

from the Compton's effect

del lamda= hc/(mc^2)*(1-cos(theta))

for maximum transfer of energy theta should be 180

del lamda= (1240 eVnm)/(511000 eV)*(2)
=0.00485 nm

E=37.7 keV=hc/lamda

=1240eVnm/lamda
lamda=0.03289 nm

lamda'=0.03289+0.00485 = 0.03774 nm

E'=hc/ lamda '

= 1240eV nm/0.03774 nm

= 32.856 keV

E=E'+Ee

Ee=E-E'=37.7-32.856= 4.84 keV

(b)

the photon scatters at an angle of pi radians or 180 degree

(c)

the electron scatters at an angle of zero radians.